OpenStudy (anonymous):

integral from 1 to 2 1/(x(lnx))

OpenStudy (anonymous):

\[\int\limits_{1}^{2} 1/(xln(x)) dx\] \[\int\limits_{1}^{2} (1/x) (1/ln(x)) dx\] \[u=\ln(x)\] \[du=(1/x)dx\] \[\int\limits_{0}^{ln(2)} (1/u) du\] \[\int\limits_{0}^{ln(2)}u ^{-1}du\] Think you got it from there?

OpenStudy (callisto):

I found something weird...